# DROP: Poison Dilution via Knowledge Distillation for Federated Learning

### v1

@@ 4. Our Defense: DROP
The first countermeasure, Agglomerative Clustering (Section 4.1), identifies and isolates malicious updates by clustering submitted models based on their pairwise Euclidean distances. This process is particularly effective against high-DPR attacks, where malicious updates deviate significantly from benign updates. The second countermeasure, Activity Monitoring (Section 4.2), tracks client behavior across training rounds. By maintaining a reputation score based on prior clustering results, this mechanism penalizes clients frequently flagged as suspicious and reduces their influence in future aggregations. This reputation-based approach ensures robustness against per round benign-to-malicious client ratio fluctuations, where the proportion of malicious clients can vary widely across rounds due to random client selection.
The third and final countermeasure, Knowledge Distillation (Section 4.3), addresses the most challenging class of attacks: stealthy, low-DPR strategies. These attacks encode adversarial objectives subtly into model updates, making them indistinguishable from benign updates and enabling them to bypass clustering schemes. To neutralize these residual adversarial signals, we use synthetic data generated by a GAN trained via logit-driven distillation to produce a cleansed version of the global model. This process ensures robustness against low-DPR attacks and renders the defense agnostic to learning configurations.
Through this cascade of countermeasures, DROP achieves the following three goals:
  1. (1)
    Resilience against aggressive, high data poisoning rates.
  2. (2)
    Adaptability to diverse malicious client ratios.
  3. (3)
    Robustness against stealthy, low data poisoning rates.
@@
4.1. Agglomerative Clustering
The first line of defense in our system is Agglomerative Clustering (AC) (Müllner, 2011), selected for its flexibility and hierarchical structure, which make it well-suited for detecting subtle deviations caused by targeted backdoor attacks. Clustering is a widely adopted countermeasure in filtering-based FL frameworks (Nguyen et al., 2022; Fung et al., 2018; Kumari et al., 2023; Krauß and Dmitrienko, 2023; Shen et al., 2016) due to its ability to identify dissimilarities within a population of model updates, often resulting from adversarial strategies that embed backdoors aggressively with a high DPR. We chose AC over methods like K-means and HDBSCAN due to its lower sensitivity to hyper-parameter selection and its dynamic merging process, which eliminates the need to pre-specify the number of clusters. Unlike K-means, which can be sensitive to initial centroid selection and outliers, AC provides more consistent results by iteratively merging clusters based on similarity. Similarly, while HDBSCAN (Campello et al., 2013) relies on density-based assumptions that require careful tuning of hyper-parameters, AC operates without such assumptions, making it more robust in diverse FL scenarios.
The use of Ward linkage (Jr., 1963) further enhances AC’s effectiveness by ensuring that intra-cluster variance is minimized during the merging process. Among the various linkage methods available for AC, Ward linkage is particularly effective for FL tasks, as it prioritizes clusters with low internal variance, which is crucial for distinguishing fine-grained and stealthy backdoors embedded within benign-looking updates. By capturing and separating subtle differences in high-dimensional weight spaces, Ward linkage ensures that adversarial updates are reliably isolated from legitimate ones.
@@ 4.1. Agglomerative Clustering
(5) $d_{\text{Ward}}(\mathcal{A},\mathcal{B})=\frac{|\mathcal{A}|\,|\mathcal{B}|}{|\mathcal{A}|+|\mathcal{B}|}\,\|\boldsymbol{\mu}_{\mathcal{A}}-\boldsymbol{\mu}_{\mathcal{B}}\|_{2}^{2},$
@@ 4.2. Activity Monitoring
This mechanism addresses a limitation of clustering-based defenses in federated learning, which assume that the number of malicious updates in a round is smaller than the number of benign ones. In real-world scenarios, however, client selection is random, and the defender has no prior knowledge of the number of malicious clients in any given round. This randomness can lead to rounds where adversarial clients outnumber benign ones, causing misidentification of the smaller cluster. Prior works often assume an upper bound on the proportion of malicious clients (typically less than 50%), but enforcing this requires low MCRs or artificially controlled client sampling to maintain a fixed benign-to-malicious client ratio across rounds (Zhang et al., 2023), limiting practical applicability. It is important to note that although the total proportion of adversarial clients in the federation (MCR) remains fixed, the actual ratio of malicious clients in any given round can fluctuate due to the random selection of clients.
@@
(6) $P\left(\mathcal{M}\geq\frac{C}{2}\right)\geq 1-\left(4\rho(1-\rho)\right)^{\frac{C}{2}}.$
For a derivation of the above, please see Appendix A. This bound implies that as $C$ increases, the likelihood of selecting a disproportionately large number of malicious clients diminishes exponentially. Conversely, when $C$ is small, the probability that malicious clients outnumber benign clients within the selected subset increases. For example, consider an FL system with $N=100$ total clients and a MCR of 40% ($\rho=0.4$). If $C=20$ clients are randomly selected in a given round, the expected number of malicious clients in the subset is $\mathbb{E}[\mathcal{M}]=\rho\cdot C=0.4\cdot 20=8$. However, with the inequality above we can see the probability of selecting more than 10 malicious clients can be non-trivial ($\approx 0.34$), potentially resulting in a subset where malicious clients constitute more than 50% of the selected participants ($M>C/2$). In fact, the bound suggests that in about a third of the FL training rounds, the malicious clients will be the majority.
To address scenarios where the number of malicious clients exceeds the benign ones in certain rounds, we propose a reputation-based mechanism. This approach tracks client behavior across rounds, penalizing suspicious clients identified by the clustering component (Section 4.1) and gradually reducing their influence over the global model. By doing so, our method dynamically adjusts to varying levels of malicious participation without relying on rigid assumptions about MCR or artificially controlling client sampling. For instance, FLIP (Zhang et al., 2023) enforces a fixed per-round MCR by ’randomly’ selecting 10 clients per round, ensuring exactly 4 adversaries and 6 benign clients, thus artificially tampering with the randomness of client selection. By incorporating a reputation system, our defense mitigates errors caused by transient imbalances where adversarial clients may temporarily outnumber benign ones. Over multiple rounds, as long as the overall MCR remains below 50%, the mechanism ensures that benign clients are correctly identified and retained more frequently than adversarial clients. This design helps prevent long-term accumulation of adversarial influence, even in rounds where clustering misclassifications may occur.
@@
Calculating Penalty Scores. We begin by initializing the penalty score $\pi(c)=0$ for every client. As training proceeds, this score is dynamically updated based on the clustering results from each round. During each training round $t$, the clustering component (Section 4.1) determines whether a client’s update is benign or suspicious by assigning the client with a penalty score:
(7) $\pi(c)=\begin{cases}\pi(c)+p,&\text{if }c\in\mathcal{C}_{s}\\ \max(0,\pi(c)-r),&\text{otherwise}\end{cases}$
@@
The server uses these scores to regulate client participation. Clients with a clean history i.e., consistently flagged as benign ($\pi(c)=0$) are deemed trustworthy and allowed to contribute to the global model, whereas clients with a record of suspicious update submissions ($\pi(c)>0$) are restricted from participating in the aggregation process. However, depending on the server administration policy, the penalty system could be even stricter. Clients that accumulate excessive penalty points due to repeated suspicious updates are permanently blacklisted from contributing in subsequent rounds: Optionally, the server could permanently ban any client whose penalty score exceeds a predefined threshold.
@@ 4.3. Knowledge Distillation
This approach leverages model stealing attacks, such as the MAZE model stealing framework (Kariyappa et al., 2021), which achieves state-of-the-art performance in extracting machine learning models under black-box conditions. The goal here is not to engage in model stealing as traditionally intended, but to clone the model without introducing any malicious behavior from compromised clients. MAZE combines two key components: query synthesis and knowledge distillation. Knowledge Distillation (Hinton et al., 2015) is a technique where a ”student” model learns to approximate the output logits of a larger ”teacher” model, enabling the transfer of knowledge while preserving essential decision boundaries. A generative adversarial network (GAN) (Radford et al., 2016) synthesizes inputs to probe a target model’s decision boundaries, and knowledge distillation transfers the target model’s behavior to a surrogate model trained on these synthetic queries. By minimizing reliance on labeled data, MAZE efficiently approximates the target model’s functionality.
In our approach, we adapt and extend the MAZE framework to take advantage of the white-box access available to the server in FL, as the server owns the global model. This adjustment eliminates the need for zeroth-order gradient estimation used in black-box settings and allows for direct backpropagation, improving both efficiency and accuracy. To ensure high-quality synthetic data, we initialize the GAN with a set of $n$ clean samples, denoted as $\mathcal{D}_{\text{clean}}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ in step 4 of Algorithm 1, which helps align the generated queries with the original data distribution. Our approach requires only a small set of clean samples—typically less or equal than the size of a single client’s dataset—making it practical and feasible for real-world FL scenarios.
@@
(8) $\mathbf{z}_{t}^{c}=f(\mathbf{x};\mathbf{w}_{t}^{c}),\quad\forall c\in\mathcal{C}_{b}.$
@@
(9) $\mathbf{\bar{z}}_{t}=\frac{1}{|\mathcal{C}_{b}|}\sum_{c\in\mathcal{C}_{b}}\mathbf{z}_{t}^{c}.$
The generator $G$ uses the ensemble logits $\mathbf{\bar{z}}_{t}$ as feedback to synthesize new queries $\mathcal{D}_{\text{synthetic}}$. These synthetic queries, denoted as $\mathbf{x}_{\text{gen}}$, are used to guide a clone network, which serves as a cleansed version of the global model. The clone network, parameterized by $\mathbf{w}^{\text{clone}}_{t}$, is trained to align its predictions with the ensemble logits. Instead of the KL divergence used in the original MAZE framework, we employ the $\ell_{1}$-loss, which provides more stable performance (Truong et al., 2021). The training objective for the clone network is defined as:
(10) $\mathcal{L}_{\text{distill}}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{synthetic}}}\left[\|\mathbf{\bar{z}}_{t}-\mathbf{z}_{t}^{\text{clone}}\|_{1}\right],$
@@
4.4. DROPlet: a Lightweight, Scalable Defense Mechanism
@@
Despite its lightweight design, DROPlet offers robust protection by leveraging clustering to detect anomalies and activity monitoring to penalize malicious clients. However, it may not be effective against stealthy, low-DPR attacks (Section 5.2.3). Nevertheless, DROPlet remains a competitive baseline for FL defense evaluations, offering significant protection while being faster and easier to deploy compared to classical methods like Multi-KRUM and Median.
5. Experiments
We conduct extensive experiments under a wide range of adversarial conditions, data distributions, and model architectures. Our evaluations focus on various learning configurations of interest, as discussed in Section 3.2. We also vary DPRs, MCRs, and use both independent and identically distributed (IID) and non-IID data distributions, ensuring a comprehensive assessment of DROP in realistic FL settings.
5.1. Setup
Datasets and Model Architecture. We evaluated our defense using three standard FL datasets: CIFAR-10, CINIC-10, and EMNIST. CIFAR-10 (Krizhevsky et al., 2009) contains 60,000 32x32 color images across 10 diverse classes. CINIC-10 (Darlow et al., 2018) is an extension of CIFAR-10, incorporating additional images from ImageNet (Deng et al., 2009) to increase dataset size and variability. It consists of 270,000 images distributed across the same 10 classes as CIFAR-10, providing a more challenging classification task with greater intra-class diversity. EMNIST (Cohen et al., 2017) extends MNIST (Deng, 2012) with 62 handwritten character and digit classes, making it suitable for FL classification tasks. We use a ResNet-18 architecture (He et al., 2015) for all experiments.
Data Distribution. We evaluate our defense on IID and across various levels of non-IID data to reflect the heterogeneity typically observed in real-world FL deployments. To control the degree of non-IID-ness, we partition the global dataset using a Dirichlet distribution $Dir(\alpha)$ (Li et al., 2021) (see Section 5.2) and use the distributional parameter $\alpha$ to vary the degree of non-IID-ness.
Federation Settings. The federation consists of 50, 100, and 150 clients for CIFAR-10, EMNIST, and CINIC-10 respectively. 50% clients are randomly selected to participate each training round for CIFAR-10 and EMNIST, while for CINIC-10 20% clients are randomly selected per training round. Unlike prior works (Blanchard et al., 2017; Zhang et al., 2023) that assume a fixed number of adversarial clients in every round $t$, our model allows for the number of adversarial clients in $\mathcal{C}_{t}$ to vary.
@@ 5.1. Setup
Undefended 87.55 87.32 87.40 85.5 96.6 98.4 74.61 74.49 94.6 98.65 89.20 89.11 98.50 97.75
FLAME 84.86 85.00 85.03 88.7 96.8 99.4 73.16 72.77 96.63 98.6 87.83 88.0 5.0 12.75
FLARE 85.06 81.08 84.41 92.5 75.8 3.5 72.18 71.09 96.71 93.02 89.07 88.92 96.00 84.50
FLIP 83.09 81.53 82.88 53.4 73.4 48.80 70.08 69.60 80.98 38.82 88.53 88.32 46.50 48.75
FLTrust 87.93 87.52 87.62 90.4 96.7 98.1 74.74 74.54 90.83 98.72 89.19 89.18 97.0 90.0
Fool’s Gold 83.85 83.76 83.92 83.0 92.6 94.3 70.05 69.96 89.33 95.29 87.85 87.80 77.0 46.50
Median 87.48 87.33 87.21 95.3 96.3 98.8 73.49 73.38 97.6 98.87 89.20 89.20 96.75 96.25
Multi-KRUM 85.96 86.00 85.87 86.0 97.5 98.6 73.57 73.58 76.35 99.74 88.70 88.58 78.75 5.0
@@
Table 3. MTA (%) and ASR (%) metrics at different Data Poisoning Rates (DPR) for the CIFAR-10, CINC-10 and EMNIST datasets. The MCR is fixed at 20%. DROP achieves low ASR for all DPR, while existing defenses fail to mitigate the attack. The lightweight variant DROPlet achieves higher MTA than DROP and is robust in all cases, except for some extremely stealthy attacks (1.25% DPR, CIFAR-10). The learning configurations are C4 for CIFAR-10 (same for CINIC-10) and C5 for EMNIST. Refer to Table 2 and Table 7 for learning configuration details.
@@
Attacks. We evaluate our defense against a targeted backdoor attack (described in Section 2.2), where adversarial clients poison inputs from a specific victim class $y_{\text{victim}}$ to embed malicious behavior in the global model. The adversary applies a trigger $\mathbf{t}$ to a fraction of victim-class inputs and flips their labels to a target class $y_{\text{target}}$. Attack success is measured by the Attack Success Rate (ASR), while maintaining high Main Task Accuracy (MTA) ensures stealthiness. We evaluate robustness by varying the DPR at 1.25%, 2.5%, and 5%, and the MCR at 20% and 40%, to simulate a range of adversarial compromise.
@@
Baselines. To benchmark the effectiveness of DROP, we compare its performance against representative FL poisoning defense mechanisms from the literature. Median Aggregation (Yin et al., 2018) is a classical robust aggregation technique that uses the coordinate-wise median of client updates to neutralize the effect of outliers. Multi-Krum (Blanchard et al., 2017) selects and aggregates the most ”trustworthy” client updates by iteratively excluding updates that are furthest from the majority, making it resilient to Byzantine failures. FLTrust (Cao et al., 2021) introduces a server-side reference model to measure the ”trustworthiness” of client updates, ensuring that only updates aligning with the reference model are included in the aggregation. FoolsGold (Fung et al., 2018) employs similarity-based clustering to mitigate the impact of poisoned updates by reducing their contribution in the aggregation process. FLAME (Nguyen et al., 2022) is a robust aggregation method designed to address both Byzantine and poisoning attacks in FL by selectively penalizing anomalous updates. FLIP (Zhang et al., 2023) estimates client-side trigger reconstruction using adversarial training combined with low-confidence refusals. Finally, FLARE (Wang et al., 2022) leverages penultimate layer representations (PLR) to evaluate the trustworthiness of model updates, assigning trust scores based on PLR discrepancies and aggregating updates weighted by these scores to defend against model poisoning attacks. For further implementation details on the baselines, please see Section C.1.
5.2. Results

### v2

@@ IV Our Defense: DROP
The first countermeasure, Agglomerative Clustering (Section IV-A), identifies and isolates malicious updates by clustering submitted models based on their pairwise Euclidean distances. This process is particularly effective against high-DPR attacks, where malicious updates deviate significantly from benign updates. The second countermeasure, Activity Monitoring (Section IV-B), tracks client behavior across training rounds. By maintaining a reputation score based on prior clustering results, this mechanism penalizes clients frequently flagged as suspicious and reduces their influence in future aggregations. This reputation-based approach ensures robustness against per round benign-to-malicious client ratio fluctuations, where the proportion of malicious clients can vary widely across rounds due to random client selection.
The third and final countermeasure, Knowledge Distillation (Section IV-C), addresses the most challenging class of attacks: stealthy, low-DPR strategies. These attacks encode adversarial objectives subtly into model updates, making them indistinguishable from benign updates and enabling them to bypass clustering schemes. To neutralize these residual adversarial signals, we use synthetic data generated by a GAN trained via logit-driven distillation to produce a cleansed version of the global model. This process ensures robustness against low-DPR attacks and renders the defense agnostic to learning configurations.
DROP is the first FL defense that leverages knowledge distillation to remove stealthy poisoning attacks. DROP’s main innovation is in its architecture that leverages these three components to counteract a wide range of adversarial strategies. In particular, through this novel set of countermeasures, DROP provides resilience against aggressive, high data poisoning rates, adaptability to diverse malicious client ratios, and robustness against stealthy, low data poisoning rates, across learning configurations.
@@
IV-A Agglomerative Clustering
The first line of defense in our system is Agglomerative Clustering (AC) [40], selected for its flexibility and hierarchical structure, which makes it well-suited for detecting subtle deviations caused by targeted backdoor attacks. Clustering is a widely adopted countermeasure in filtering-based FL frameworks [24, 21, 41, 26, 22] due to its ability to identify dissimilarities within model updates, often resulting from adversarial strategies that embed backdoors aggressively with a high DPR. We chose AC over methods like K-means and HDBSCAN due to its lower sensitivity to hyper-parameter selection and its dynamic merging process, which eliminates the need to pre-specify the number of clusters. Unlike K-means, which can be sensitive to initial centroid selection and outliers, AC provides more consistent results by iteratively merging clusters based on similarity. Similarly, while HDBSCAN [42] relies on density-based assumptions that require careful tuning of hyper-parameters, AC operates without such assumptions, making it more robust in diverse FL scenarios.
The use of Ward linkage [43] further enhances AC’s effectiveness by ensuring that intra-cluster variance is minimized during the merging process. Among the various linkage methods available for AC, Ward linkage is particularly effective for FL tasks, as it prioritizes clusters with low internal variance, which is crucial for distinguishing fine-grained and stealthy backdoors embedded within benign-looking updates. By capturing and separating subtle differences in high-dimensional weight spaces, Ward linkage ensures that adversarial updates are reliably isolated from legitimate ones.
@@ IV-A Agglomerative Clustering
$d_{\text{Ward}}(\mathcal{A},\mathcal{B})=\frac{|\mathcal{A}|\,|\mathcal{B}|}{|\mathcal{A}|+|\mathcal{B}|}\,\|\boldsymbol{\mu}_{\mathcal{A}}-\boldsymbol{\mu}_{\mathcal{B}}\|_{2}^{2},$ (5)
@@ IV-B Activity Monitoring
This mechanism addresses a limitation of clustering-based defenses in federated learning, which assume that the number of malicious updates in a round is smaller than the number of benign ones. In real-world scenarios, however, client selection is random, and the defender has no prior knowledge of the number of malicious clients in any given round. This randomness can lead to rounds where adversarial clients outnumber benign ones, causing misidentification of the smaller cluster. Prior works often assume an upper bound on the proportion of malicious clients (typically less than 50%), but enforcing this requires low MCRs or artificially controlled client sampling to maintain a fixed benign-to-malicious client ratio across rounds [25], limiting practical applicability. It is important to note that although the total proportion of adversarial clients in the federation (MCR) remains fixed, the actual ratio of malicious clients in any given round can fluctuate due to the random selection of clients.
@@
$P\left(\mathcal{M}\geq\frac{C}{2}\right)\geq 1-\left(4\rho(1-\rho)\right)^{\frac{C}{2}}.$ (6)
For a derivation of the above, please see Section -A. This bound implies that as $C$ increases, the likelihood of selecting a disproportionately large number of malicious clients diminishes exponentially. Conversely, when $C$ is small, the probability that malicious clients outnumber benign clients within the selected subset increases. For example, consider an FL system with $N=100$ total clients and a MCR of 40% ($\rho=0.4$). If $C=20$ clients are randomly selected in a given round, the expected number of malicious clients in the subset is $\mathbb{E}[\mathcal{M}]=\rho\cdot C=0.4\cdot 20=8$. However, with the inequality above we can see the probability of selecting more than 10 malicious clients can be non-trivial ($\approx 0.34$), potentially resulting in a subset where malicious clients constitute more than 50% of the selected participants ($M>C/2$). In fact, the bound suggests that in about a third of the FL training rounds, the malicious clients will be the majority.
To address scenarios where the number of malicious clients exceeds the benign ones in certain rounds, we propose a reputation-based mechanism. This approach tracks client behavior across rounds, penalizing suspicious clients identified by the clustering component (Section IV-A) and gradually reducing their influence over the global model. By doing so, our method dynamically adjusts to varying levels of malicious participation without relying on rigid assumptions about MCR or artificially controlling client sampling. For instance, FLIP [25] enforces a fixed per-round MCR by ’randomly’ selecting 10 clients per round, ensuring exactly 4 adversaries and 6 benign clients, thus artificially tampering with the randomness of client selection. By incorporating a reputation system, our defense mitigates errors caused by transient imbalances where adversarial clients may temporarily outnumber benign ones. Over multiple rounds, as long as the overall MCR remains below 50%, the mechanism ensures that benign clients are correctly identified and retained more frequently than adversarial clients. This design helps prevent long-term accumulation of adversarial influence, even in rounds where clustering misclassifications may occur.
@@
Calculating Penalty Scores. We begin by initializing the penalty score $\pi(c)=0$ for every client. As training proceeds, this score is dynamically updated based on the clustering results from each round. During each training round $t$, the clustering component (Section IV-A) determines whether a client’s update is benign or suspicious by assigning the client with a penalty score:
$\pi(c)=\begin{cases}\pi(c)+p,&\text{if }c\in\mathcal{C}_{s}\\ \max(0,\pi(c)-r),&\text{otherwise}\end{cases}$ (7)
@@
The server uses these scores to regulate client participation. Clients with a clean history i.e., consistently flagged as benign ($\pi(c)=0$) are deemed trustworthy and allowed to contribute to the global model, whereas clients with a record of suspicious update submissions ($\pi(c)>0$) are restricted from participating in the aggregation process. However, depending on the server administration policy, the penalty system could be even stricter. Clients that accumulate excessive penalty points due to repeated suspicious updates are permanently blacklisted from contributing in subsequent rounds:
Optionally, the server could permanently ban any client whose penalty score exceeds a predefined threshold.
@@ IV-C Knowledge Distillation
This approach leverages model stealing attacks, such as the MAZE model stealing framework [44], which achieves state-of-the-art performance in extracting machine learning models under black-box conditions. The goal here is not to engage in model stealing as traditionally intended, but to clone the model without introducing any malicious behavior from compromised clients. MAZE combines two key components: query synthesis and knowledge distillation. Knowledge Distillation [45] is a technique where a ”student” model learns to approximate the output logits of a larger ”teacher” model, enabling the transfer of knowledge while preserving essential decision boundaries. A generative adversarial network (GAN) [46] synthesizes inputs to probe a target model’s decision boundaries, and knowledge distillation transfers the target model’s behavior to a surrogate model trained on these synthetic queries. By minimizing reliance on labeled data, MAZE efficiently approximates the target model’s functionality.
In our approach, we adapt and extend the MAZE framework to take advantage of the white-box access available to the server in FL, as the server owns the global model. This adjustment eliminates the need for zeroth-order gradient estimation used in black-box settings and allows for direct backpropagation, improving both efficiency and accuracy. To ensure high-quality synthetic data, we initialize the GAN with a set of $n$ clean samples, denoted as $\mathcal{D}_{\text{clean}}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ in step 4 of Algorithm 1, which helps align the generated queries with the original data distribution. These samples are drawn from the same data distribution as that of the federation clients and serve as a lightweight initialization prior for the generator. Our approach requires only a small set of clean samples—typically less than or equal to the size of a single client’s dataset—ensuring that the additional data requirement remains realistic and practical for deployment in real-world FL scenarios.
@@
$\mathbf{z}_{t}^{c}=f(\mathbf{x};\mathbf{w}_{t}^{c}),\quad\forall c\in\mathcal{C}_{b}.$ (8)
@@
$\mathbf{\bar{z}}_{t}=\frac{1}{|\mathcal{C}_{b}|}\sum_{c\in\mathcal{C}_{b}}\mathbf{z}_{t}^{c}.$ (9)
The generator $G$ uses the ensemble logits $\mathbf{\bar{z}}_{t}$ as feedback to synthesize new queries $\mathcal{D}_{\text{synthetic}}$. These synthetic queries, denoted as $\mathbf{x}_{\text{gen}}$, are used to guide a clone network, which serves as a cleansed version of the global model. The clone network, parameterized by $\mathbf{w}^{\text{clone}}_{t}$, is trained to align its predictions with the ensemble logits. Instead of the KL divergence used in the original MAZE framework, we employ the $\ell_{1}$-loss, which provides more stable performance [47]. The training objective for the clone network is defined as:
$\mathcal{L}_{\text{distill}}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{synthetic}}}\left[\|\mathbf{\bar{z}}_{t}-\mathbf{z}_{t}^{\text{clone}}\|_{1}\right],$ (10)
@@
IV-D DROPlet: a Lightweight, Scalable Defense Mechanism
@@
Despite its lightweight design, DROPlet offers robust protection by leveraging clustering to detect anomalies and activity monitoring to penalize malicious clients. However, it may not be effective against stealthy, low-DPR attacks. Nevertheless, DROPlet remains a competitive baseline for FL defense evaluations, offering significant protection while being faster and easier to deploy compared to classical methods like Multi-KRUM and Median.
V Experiments
We conduct extensive experiments under a wide range of adversarial conditions and data distributions. Our evaluations focus on various learning configurations of interest, as discussed in Section III-B. We also vary attack strategies, DPRs, MCRs, and use both independent and identically distributed (IID) and non-IID data distributions, ensuring a comprehensive assessment of DROP in realistic FL settings.
Datasets and Model Architecture. We evaluated our defense using three standard FL datasets: CIFAR-10, CINIC-10, and EMNIST. CIFAR-10 [48] contains 60,000 32x32 color images across 10 diverse classes. CINIC-10 [49] is an extension of CIFAR-10, incorporating additional images from ImageNet [50] to increase dataset size and variability. It consists of 270,000 images distributed across the same 10 classes as CIFAR-10, providing a more challenging classification task with greater intra-class diversity. EMNIST [51] extends MNIST [52] with 62 handwritten character and digit classes, making it suitable for FL classification tasks. We use a ResNet-18 architecture [53] for all experiments.
Data Distribution. We evaluate our defense on IID and across various levels of non-IID data to reflect the heterogeneity typically observed in real-world FL deployments. To control the degree of non-IID-ness, we partition the global dataset using a Dirichlet distribution $Dir(\alpha)$ [54] (see Section V-D) and use the distributional parameter $\alpha$ to vary the degree of non-IID-ness.
Federation Settings. The federation consists of 50, 100, and 150 clients for CIFAR-10, EMNIST, and CINIC-10 respectively. 50% clients are randomly selected to participate each training round for CIFAR-10 and EMNIST while for CINIC-10, 20% clients are randomly selected per training round. Unlike prior works [19, 25] that assume a fixed number of adversarial clients in every round $t$, our model allows for the number of adversarial clients in $\mathcal{C}_{t}$ to vary.
@@ V Experiments
DPR (%) NA 1.25 2.5 5 1.25 2.5 5 NA 1.25 2.5 1.25 2.5 NA 1.25 2.5 1.25 2.5
Undefended 87.56 87.55 87.32 87.40 85.5 96.6 98.4 74.71 74.61 74.49 94.6 98.65 89.01 89.20 89.11 98.50 97.75
FLAME 85.12 84.86 85.00 85.03 88.7 96.8 99.4 72.93 73.16 72.77 96.63 98.6 88.02 87.83 88.0 5.0 12.75
FLARE 85.29 85.06 81.08 84.41 92.5 75.8 3.5 71.07 72.18 71.09 96.71 93.02 88.82 89.07 88.92 96.00 84.50
FLIP 81.25 83.09 81.53 82.88 53.4 73.4 48.80 67.73 70.08 69.60 80.98 38.82 88.31 88.53 88.32 46.50 48.75
FLTrust 87.68 87.93 87.52 87.62 90.4 96.7 98.1 74.46 74.74 74.54 90.83 98.72 89.22 89.19 89.18 97.0 90.0
Fool’s Gold 83.59 83.85 83.76 83.92 83.0 92.6 94.3 70.25 70.05 69.96 89.33 95.29 87.92 87.85 87.80 77.0 46.50
Median 87.07 87.48 87.33 87.21 95.3 96.3 98.8 73.29 73.49 73.38 97.6 98.87 88.94 89.20 89.20 96.75 96.25
Multi-KRUM 85.66 85.96 86.00 85.87 86.0 97.5 98.6 73.61 73.57 73.58 76.35 99.74 88.62 88.70 88.58 78.75 5.0
DROPlet (ours) 87.33 87.49 87.05 87.20 91.3 0.1 0.9 74.62 74.05 74.27 15.96 0.48 89.15 88.98 88.96 1.50 1.00
DROP (ours) 76.07 76.05 76.53 75.18 1.3 1.1 0.5 65.4 65.65 65.72 1.22 1.44 89.32 88.62 89.13 0.75 1.00
TABLE III: MTA (%) and ASR (%) metrics under no attack (NA) and at different Data Poisoning Rates (DPR) for the CIFAR-10, CINIC-10 and EMNIST datasets. The MCR is fixed at 20%. DROP achieves low ASR for all DPR, while existing defenses fail to mitigate the attack. The lightweight variant DROPlet achieves higher MTA than DROP and is robust in all cases, except for some extremely stealthy attacks (1.25% DPR, CIFAR-10). The learning configurations are C4 for CIFAR-10 (same for CINIC-10) and C5 for EMNIST. Refer to Table II and Table VIII for learning configuration details.
Attack Defense CIFAR-10 CINIC-10 EMNIST
MTA (%) ASR (%) MTA (%) ASR (%) MTA (%) ASR (%)
@@
BadNets Undefended 87.55 87.32 87.40 85.5 96.6 98.4 74.61 74.49 94.6 98.65 89.20 89.11 98.50 97.75
FLIP 83.09 81.53 82.88 53.4 73.4 48.80 70.08 69.60 38.82 88.53 88.53 88.32 46.50 48.75
@@
Neurotoxin Undefended 87.64 87.27 87.09 92.20 93.20 98.2 74.24 74.29 97.50 98.54 89.16 89.17 98.0 97.75
FLIP 82.76 83.18 82.04 53.0 84.40 73.4 69.83 69.87 24.52 77.27 88.52 88.49 41.25 10.75
DROPlet (ours) 87.10 87.36 87.08 88.80 0.50 0.7 74.02 74.37 97.0 0.69 89.03 88.95 2.25 1.50
DROP (ours) 77.06 76.66 75.03 1.20 1.60 1.1 65.78 65.48 1.20 1.02 89.36 89.45 0.75 1.0
Chameleon Undefended 87.06 87.34 87.43 100.0 100.0 100.0 74.76 74.34 99.96 99.91 89.02 88.98 99.25 99.5
FLIP 83.29 83.41 83.64 96.8 97.1 98.0 70.31 70.01 88.91 97.97 88.48 88.65 77.25 25.25
DROPlet (ours) 87.72 87.1 87.33 99.8 99.9 0.4 74.11 74.4 0.79 0.72 88.98 88.95 1.25 1.25
DROP (ours) 75.71 75.37 76.14 1.0 1.2 1.8 65.98 66.06 1.48 1.87 89.32 89.43 1.0 1.0
@@
TABLE IV: MTA (%) and ASR (%) metrics for different attack strategies and at different Data Poisoning Rates (DPR) for the CIFAR-10, CINIC-10, and EMNIST datasets. The MCR is fixed at 20%. DROP achieves low ASR for all DPR values across all attacks. DROPlet maintains higher MTA than DROP and remains robust in most cases, though it is occasionally affected by Chameleon’s adaptive poisoning (e.g., 2.5% DPR, CIFAR-10). The learning configurations are C4 for CIFAR-10 (same for CINIC-10) and C5 for EMNIST. Refer to Table II and Table VIII for learning configuration details.
@@
Attacks. We evaluate our defense against a targeted backdoor attack (described in Section II-B), where adversarial clients poison inputs from a specific victim class $y_{\text{victim}}$ to embed malicious behavior in the global model. We evaluate over BadNets [6], as well as two state-of-the-art adaptive variants: Neurotoxin [55] and Chameleon [56]. Neurotoxin improves durability by constraining poisoned updates to model parameters that typically change little during training, embedding a backdoor that persists even after the adversary drops out. Chameleon uses contrastive learning to adapt poisoned inputs to the feature space of benign peer samples, aligning with facilitator images and evading detection. Both attacks are stealthy—causing minimal degradation in clean accuracy—and durable, maintaining high attack success rates across many training rounds. The adversary applies a trigger $\mathbf{t}$ to a fraction of victim-class inputs and flips their labels to a target class $y_{\text{target}}$. Attack success is measured by the Attack Success Rate (ASR), while maintaining high Main Task Accuracy (MTA) ensures stealthiness. We evaluate robustness by varying the DPR at 1.25%, 2.5%, and 5%, and the MCR at 20% and 40%, to simulate a range of adversarial compromise.
@@
Baselines. To benchmark the effectiveness of DROP, we compare its performance against representative FL poisoning defense mechanisms from the literature. Median Aggregation [18] is a classical robust aggregation technique that uses the coordinate-wise median of client updates to neutralize the effect of outliers. Multi-Krum [19] selects and aggregates the most “trustworthy” client updates by iteratively excluding updates that are furthest from the majority, making it resilient to Byzantine failures. FLTrust [20] introduces a server-side reference model to measure the “trustworthiness” of client updates, ensuring that only updates aligning with the reference model are included in the aggregation. FoolsGold [21] employs similarity-based clustering to mitigate the impact of poisoned updates by reducing their contribution in the aggregation process. FLAME [24] is a robust aggregation method designed to address both Byzantine and poisoning attacks in FL by selectively penalizing anomalous updates. FLIP [25] estimates client-side trigger reconstruction using adversarial training combined with low-confidence refusals. Finally, FLARE [23] leverages penultimate layer representations (PLR) to evaluate the trustworthiness of model updates, assigning trust scores based on PLR discrepancies and aggregating updates weighted by these scores to defend against model poisoning attacks. For further implementation details on the baselines, please see Section -C1.